# Embodied LLM Agents Learn to Cooperate in Organized Teams

### v1

@@ 4.1 A Designated Leader Enhances Performance
We first studied the effect of organizational structures and leadership on LLM agents. For benchmarking, we experimented with disorganized LLM agents without providing any organization prompt. In this case, agents still communicate with one another and work to complete the overall task. However, we discovered frequent occasions where agents send redundant, repetitive messages and interfere with one another. See Figure 1 for an illustration and see Appendix D for more examples. Numeric metrics are reported in Appendix Table 1.
@@
The proposed multi-LLM-agent architecture is also human-friendly to support human-AI collaboration. In the experiment, we ask a human player to replace the leader in the team of 3 GPT-4 agents. We recruit three human players to conduct the experiments. Figure 4(a, b) demonstrates that human leadership achieved better task completion time and improved communication efficiency compared with GPT-4 as the leader. Please find more examples of dialogues between the human leader and LLM agents in Appendix F.
@@ 4.2 Leadership and Open Communication Matters
LLM agents have different levels of leadership. In the team with a mixture of GPT-4 and GPT-3.5-turbo agents, appointing GPT-4 as the leader increases the team efficiency higher than if GPT-3.5-turbo is the leader (Figure 4(c,d)). We ran this experiment on teams of three agents and five agents, respectively. In both scenarios, the task completion time and communication cost are reduced when GPT-4 acts as the leader. This finding implies different levels of leadership between these LLMs.
We also observed that encouraging constructive feedback to the leader agent helped performance. Motivated by successful human organizations, we tried to promote open communications among LLM agents by adding an additional prompt that "If the leader’s instructions are not right, you can correct the leader". Figure 4(c, d) illustrates the results. Interestingly, this modification improves the team’s overall efficiency and reduces the time to task completion when the team is made up of 3$\times$×GPT-4 ($t(38)=0.87,p=0.14$). In contrast, the same modification lowers the team efficiency when GPT-3.5-turbo agents try to correct the leader ($t(38)=0.27,p=0.40$). In both experiments, the communication cost increases. We present more details about these behaviors in Appendix G.
@@ 4.3 Emergence of Cooperative Behaviors
We delved into the behaviors of LLM agents in an organized team to investigate how organization prompts influence agents’ communication and decisions. Analysis of their dialogue history revealed that agents demonstrated a variety of cooperative behaviors, such as reporting, correction, task allocation, and asking for help (see Figure 6 for an example dialogue).
One may argue that these types of behaviors could also emerge due to the nature of LLMs, even without a pre-specified team structure. Thus we performed a quantitative analysis to study the impact of an organization prompt on these behaviors. We followed a three-step process:
  • (1)
    We defined three major categories of human cooperative behaviors: (i) Information sharing: agents influence others by offering new information, either actively or by being asked. Reporting to the leader, sharing new observations, and answering questions belong to this category. (ii) Leadership & assistance: agents, especially the leader if there is one, can influence others by changing their plans. The behaviors include task allocation, correction, and asking for help. (iii) Request for guidance: agents actively request new information or plans for their own decision making.
  • (2)
    We developed a standalone prompt-based GPT-4- classifier to analyze each piece of dialogue. The classifier decides whether to label the dialogue with any subset of the aforementioned labels. The classifier has an accuracy of 91.67% when tested on 20 human-labeled dialogue samples with 60 labels (see Appendix A for the prompt and Appendix H for the test samples).
  • (3)
    We use the classifier to label messages generated by our teams of agents, and report the percentages of messages with cooperative behaviors. Note that a message may have multiple labels.
Figure 7 reports the results and illustrates the behavior patterns for different LLM agents. The results support several observations. Even in a disorganized team, LLM agents love to tell others what to do. Leadership & assistance accounts for around $>$ 50% of all the behaviors (Figure 7(a)). However, other than telling others what to do, agents in the disorganized team do not show much cooperative behaviours, for example they would request for guidance in $<10\%$ of the dialogues.
In contrast, when the team has a hierarchical organization the lead LLM agent would presume a dominant role and give orders to others (amount to $>60\%$ of their communication), while other members tend to follow and give fewer orders compared with the disorganized case. (Figure 7(b, c)). In such a team, agents tend to share and ask for more information, especially for the follower agents in the team.
@@ 4.4 Novel Organizational Structures
Having evaluated the merits of different kinds of structures, we let the LLM propose novel organizational structures and iteratively refine the organization prompts using the Criticize-Reflect method discussed in Section 3 (see also Figure 3).
Figure 8(a) visualizes the reflection process. The system was initialized with a basic organization prompt, i.e., “Agent_1 as the leader to coordinate the task". As the Reflection process moves forward, the Coordinator generates a sequence of evolving organization prompts, picking up key words like “hierarchical" and “dynamic" that imply more complex team structures.
@@
The Critic analyzes the records of action and dialogue, and performance metrics from the most recent episode. It provides evaluation for the full team’s trajectory, feedback to individual agents and their rankings. See below for an example of the Critic output:
@@
Figure 8: The reflection and improvement process for finding novel organizational structures. (a) The experiment was done using the 1$\times$×GPT-4+2$\times$×GPT-3.5-turbo team. The organization prompt evolves during the iterations, and takes on additional keywords such as "central", "hierarchical", and "dynamic". (b) The confidence intervals are calculated over 20 seeds.
Figure 9: Communication patterns and the corresponding organizational prompts for different team structures. (a) Team without organization prompts. (b) Team with a leader. (c) A team in the chain structure. (d) A dual-leader team. (e) A team with a dynamic leadership. (c, d, e) are proposed by LLM via Reflection. Red-robot nodes mark the lead agents, and other nodes are the followers. Edges mark the accumulated communication cost between the two nodes (darker edge means higher token costs).
@@ 5 Conclusion
We developed a novel multi-LLM-agent architecture to facilitate communication and organize the embodied agent teams for enhanced cooperation. Moreover, we proposed the Criticize-Reflect architecture based on LLMs to generate more efficient organizational prompts. Extensive experiments with various group settings and organizational structures demonstrate that a hierarchically-organized team with a designated/elected leader has superior team efficiency, which can be further improved by Criticize-Reflect.
The current work is performed in a single environment and lacks human evaluation. Future work shall expand the evaluation to a broader set of environments, allowing human evaluation and more complex human-AI collaboration.
Impact Statement
@@ Appendix A Prompt Templates
Appendix B Additional Results
B.1 Complete list of basic experimental results
@@
B.2 Across Task Generalizability
@@
Appendix C Emergent Cooperative Behaviors in an Organization
@@
Appendix D Ineffective Communication
@@
Appendix E Examples of Election
@@
Appendix F Examples of Human-AI Collaboration
@@ Appendix G Examples of Correction
Appendix H Examples of Cooperative Behaviors Classification by Humans and GPT-4
@@
Appendix I Examples of New Prompts after Reflection

### v2

@@ 4.1 A Designated Leader Enhances Performance
We first studied the effect of organizational structures and leadership on LLM agents. For benchmarking, we experimented with disorganized LLM agents without providing any organizational prompt. In this case, agents still communicate with one another and work to complete the overall task. However, we discovered frequent occasions where agents send redundant, repetitive messages and interfere with one another. See Figure 1 for an illustration and see Appendix E for more examples. Numeric metrics are reported in Appendix Table 1.
@@
The proposed multi-LLM-agent architecture is also human-friendly to support human-AI collaboration. In the experiment, we ask a human player to replace the leader in the team of 3 GPT-4 agents. We recruit three human players to conduct the experiments. Figure 4(a, b) demonstrates that human leadership achieved better task completion time and improved communication efficiency compared with GPT-4 as the leader. Please find more examples of dialogues between the human leader and LLM agents in Appendix F.2.
@@ 4.2 Leadership and Open Communication Matters
LLM agents have different levels of leadership. In the team with a mixture of GPT-4 and GPT-3.5-turbo agents, appointing GPT-4 as the leader increases the team efficiency higher than if GPT-3.5-turbo is the leader (Figure 4(c,d), Appendix F.4). We ran this experiment on teams of three agents and five agents, respectively. In both scenarios, the task completion time and communication cost are reduced when GPT-4 acts as the leader. This finding implies different levels of leadership between these LLMs.
We also observed that encouraging constructive feedback to the leader agent helped performance. Motivated by successful human organizations, we tried to promote open communications among LLM agents by adding an additional prompt that "If the leader’s instructions are not right, you can correct the leader". Figure 4(c, d) illustrates the results. Interestingly, this modification improves the team’s overall efficiency and reduces the time to task completion when the team is made up of 3$\times$×GPT-4 ($t(38)=0.87,p=0.14$). In contrast, the same modification lowers the team efficiency when GPT-3.5-turbo agents try to correct the leader ($t(38)=0.27,p=0.40$). In both experiments, the communication cost increases. We present more details about these behaviors in Appendix F.3.
@@ 4.3 Emergence of Cooperative Behaviors
We delved into the behaviors of LLM agents in an organized team to investigate how organizational prompts influence agents’ communication and decisions. Analysis of their dialogue history revealed that agents demonstrated a variety of cooperative behaviors, such as reporting, correction, task allocation, and asking for help (see Figure 6 for an example dialogue).
One may argue that these types of behaviors could also emerge due to the nature of LLMs, even without a pre-specified team structure. Thus we performed a quantitative analysis to study the impact of an organizational prompt on these behaviors. We followed a three-step process:
  • (1)
    We defined three major categories of human cooperative behaviors: (i) Information sharing: agents influence others by offering new information, either actively or by being asked. Reporting to the leader, sharing new observations, and answering questions belong to this category. (ii) Leadership & assistance: agents, especially the leader if there is one, can influence others by changing their plans. The behaviors include task allocation, correction, and asking for help. (iii) Request for guidance: agents actively request new information or plans for their own decision-making.
  • (2)
    We developed a standalone prompt-based GPT-4-classifier to analyze each piece of dialogue. The classifier decides whether to label the dialogue with any subset of the aforementioned labels. The classifier has an accuracy of 91.67% when tested on 20 human-labeled dialogue samples with 60 labels (see Appendix A for the prompt and Appendix G for the test samples).
  • (3)
    We use the classifier to label messages generated by the agents and report the percentages of messages with cooperative behaviors. Note that one message may have multiple labels.
Figure 7 reports the results and illustrates the behavior patterns for different LLM agents. The results support several observations. Even in a disorganized team, LLM agents love to tell others what to do. Leadership & assistance accounts for around $>$ 50% of all the behavior (Figure 7(a)). However, other than telling others what to do, agents in the disorganized team do not show much cooperative behavior, for example, they would request for guidance in $<10\%$ of the dialogues.
In contrast, when the team has a hierarchical organization, the lead LLM agent would presume a dominant role and give orders to others (amount to $>60\%$ of their communication), while other members tend to follow and give fewer orders compared with the disorganized case. (Figure 7(b, c)). In such a team, agents tend to share and ask for more information, especially for the follower agents in the team. But still, the agents may fail to cooperate well, such as being lazy and confused about numbers, please see examples in Appendix F.6.
@@ 4.4 Novel Organizational Structures
Having evaluated the merits of different kinds of structures, we let the LLMs propose novel organizational structures and iteratively refine the organizational prompts using the Criticize-Reflect method discussed in Section 3 (see also Figure 3).
Figure 8(a) visualizes the reflection process. The system was initialized with a basic organizational prompt, i.e., "Agent_1 as the leader to coordinate the task". As the Reflection process moves forward, the Coordinator generates a sequence of evolving organizational prompts, picking up key words like "hierarchical" and "dynamic" that imply more complex team structures.
@@
The Critic analyzes the records of action and dialogue, and performance metrics from the most recent episode. It provides evaluation for the full team’s trajectory, feedback to individual agents and their rankings. See the example of the Critic outputs in Appendix B.1.
@@
Figure 8: The reflection and improvement process for finding novel organizational structures. (a) The experiment was done using the 1$\times$×GPT-4+2$\times$×GPT-3.5-turbo team. The organizational prompt evolves during the iterations, and takes on additional keywords such as "central", "hierarchical", and "dynamic". (b) The confidence intervals are calculated over 20 seeds.
Figure 9: Communication patterns and the corresponding organizational prompts. (a) Team without organizational prompts. (b) Team with a leader. (c) A team in the chain structure. (d) A dual-leader team. (e) A team with a dynamic leadership. (c, d, e) are proposed by Criticize-Reflect. Red-robot nodes mark the lead agents, and other nodes are the followers. Edges mark the accumulated communication cost between the two nodes (darker edge means higher token cost).
@@ 5 Conclusion
We develop a novel multi-LLM-agent architecture to facilitate communication and organize the embodied agent teams for enhanced cooperation. Moreover, we propose the Criticize-Reflect framework based on LLMs to generate more efficient organizational prompts. Extensive experiments with various group settings and organizational structures demonstrate that a hierarchically-organized team with a designated/elected leader has superior team efficiency, which can be further improved by Criticize-Reflect.
The current work is performed in a single environment and lacks human evaluation. Future work shall extend to a broader set of environments, allowing human evaluation. As VirtualHome cannot hold hundreds of agents, future work can also explore larger organizations in other environments.
@@ Appendix A Prompt Templates
Appendix B Techinical Details
B.1 Details of the Critic
The Critic offers assessments of several episodes with different organizations for the Coordinator to improve the organizational prompt. The Critic does not directly influence the specific agent’s behaviors, but instead, the Critic provides insights into organization design to influence the team’s performance.
As included in the Critic’s prompt in Appendix A, the Critic will sequentially output the thoughts of this episode’s trajectories, then the summary and problems for each agent in this episode, and finally the leadership ranking of the agents. The Critic will rank the agents according to key factors of leadership: communication skills, conflict resolution skills, flexibility, and strategy. Note that we do not ask the Critic to score the agents because the scoring criteria could vary for different episodes, making the scores not comparable.
An example output of the Critic is provided as follows:
In the trajectory evaluation, the Critic compresses the trajectories with key steps and behaviors. Then the Critic gives the ranking where Agent_3 has the best leadership. Together with similar evaluations of other episodes, the Coordinator will redesign the organizational prompts, for example, Agent_3 now has more possibilities to be chosen as the leader in this case.
B.2 Details of the Coordinator
In this paper, we define organizational structure as the dynamics of information exchange among the LLM agents. Specifically, when the Coordinator generates a new organizational prompt, it contains three parts - topology, role assignment, and rules.
Here, "topology" is the type of the organization’s topology, such as decentralized, centralized with one specific leader, or pyramid. The topology can be visualized as shown in Figure 9.
"Role assignment" is the description of each agent’s duty and whether she is a leader or not. Multiple leaders with different roles are also allowed. For example, in Figure 8, the generated new prompt is “Agent_1 will act as the central coordinator, Agent_2 will execute tasks with updates only upon task completion or if issues arise, and Agent_3 will operate in a support role, assisting when called upon and avoiding repetitive queries”, giving each agent a different role.
"Rules" are the additional guidance to the agents’ behaviors, for instance, sentences like “If the leader’s instructions are not right, you can correct the leader” can be added to the new prompt.
Appendix C Additional Results
C.1 Complete list of basic experimental results
@@
C.2 Scaling up the team size
We conduct experiments with 3, 5, 7, and 9 agents to scale up the team size of 3 3$\times$×GPT-3.5-turbo agents, and observe that the communication costs increased in a nearly linear way, which suggests that our approach will not have dimension explosion when scaling up. In addition, the time to complete the task does not always improve with more agents. The performance of 9 agents (60.67$\pm$15.06) is worse than that of 7 agents (43.00$\pm$2.16), as the apartment may be too crowded to hold 9 agents.
Table 2: Performance for different team sizes.
Group setting Organization instruction Time Communication cost 3$\times$×GPT-3.5-turbo Agent 1 is the leader to coordinate the task. 92.90$\pm$14.70 59.87$\pm$6.33 5$\times$×GPT-3.5-turbo Agent 1 is the leader to coordinate the task. 80.00$\pm$20.51 132.01$\pm$5.76 7$\times$×GPT-3.5-turbo Agent 1 is the leader to coordinate the task. 43.00$\pm$2.16 233.40$\pm$70.96 9$\times$×GPT-3.5-turbo Agent 1 is the leader to coordinate the task. 60.67$\pm$15.06 296.55$\pm$65.17
C.3 Across Task Generalizability
@@
Appendix D Emergent Cooperative Behaviors in an Organization
@@
Appendix E Ineffective Communication
@@
Appendix F Examples of dialogues
F.1 Examples of Election
@@
F.2 Examples of Human-AI Collaboration
@@ F.3 Examples of Correction
F.4 Examples of Leadership Comparison
We provide more examples to compare the leadership between GPT-4 and GPT-3.5-turbo.
Figure 14: Comparison of the leadership between GPT-4 and GPT-3.5-turbo. Compared with GPT-3.5-turbo, GPT-4’s instructions are more specific, clear, and holistic.
F.5 Examples of Scaling Up
When scaling up the number of agents, the agents can emerge with more organizational structures. For example, a team of nine agents forms a pyramid structure.
Figure 15: The pyramid structure in a team of nine agents. Agent_1 is the primary leader and Agent_2 and Agent_3 are designated as vice leaders in the prompt.
F.6 Examples of Failure Cases
Though LLM agents show great capabilities to cooperate and make decisions, there are still some failure cases shown in Figure 16, such as being lazy and incorrect reasoning over the number of objects. There are also failure cases in some specific scenarios, for example, electing the leader based on hallucinations in Appendix F.1.
Figure 16: Examples of failure cases. The first case is being lazy. Instead of completing the subtask ordered by the leader Agent_1, Agent_2 directly repeats the order to Agent_3. However, this order is conflicted with the one Agent_3 received from the leader. The second case is confusion about numbers. The task demands placing two wines on the table and the team only gets one of them till now. When the leader Agent_1 asks Agent_2 to check the remaining wine, Agent_2 tries to correct the leader due to the hallucination that the team does not need an additional wine. The agents are GPT-3.5-turbo in both cases.
Appendix G Examples of Cooperative Behaviors Classification by Humans and GPT-4
@@
Appendix H Examples of New Prompts after Reflection
@@
Appendix I Broader Impacts
On the other hand, as our ability to bound and evaluate LLMs’ behaviors remains immature, when applied to human-LLM cooperative tasks, we still need to rely on some mandatory termination measures (such as human approval for high-stakes actions) instead of instructions in natural language only.